# Tripartite entanglement in $H\to ZZ,WW$ decays

### v1

@@ 1 Introduction
High-energy physics experiments offer a new playground to test quantum entanglement at the energy frontier, with proposals to test the spin entanglement of pairs of top quarks [1, 2, 3, 4, 5, 6, 7, 8], muons [9], $\tau$ leptons [10], weak bosons [11, 12, 13, 14, 15, 16, 17], as well as particles of different spin [18, 19, 20]. It is well known that, in addition to spin, particle production and decay involves orbital angular momentum (hereafter referred to as $L$). Therefore, in the aforementioned examples not only the spins of the two particles $S_{1}$, $S_{2}$ may be entangled, but they can also be entangled with $L$, as a consequence of total angular momentum conservation. An obvious step further is then to test the tripartite entanglement between $S_{1}$, $S_{2}$ and $L$ in these processes. The Hilbert space for $L$, in which the eigenstates $|lm\rangle$, with $l=0,\dots\infty$ and $m=-l,\dots,l$ form an orthonormal basis, has infinite dimension in general. However, in decay processes the conservation of total angular momentum implies that only a finite-dimensional subspace $\mathcal{H}_{L}$ is relevant. For example, in Higgs boson decays to weak boson pairs $VV$, $V=W,Z$, we have $l\leq 2$, and $\mathcal{H}_{L}$ is 9-dimensional. Likewise, in $H\to\tau^{+}\tau^{-}$, $\mathcal{H}_{L}$ has dimension four, with $l\leq 1$.
@@ 2 Higgs decay amplitudes
The angles $(\theta,\phi)$ correspond to the first decay product (with helicity $\lambda_{1}$). In the case of a scalar decay, $D^{0}_{00}=1$, and therefore the helicities are equal. Furthermore, for a scalar decay into vector bosons, the helicity amplitudes are parameterised by three quantities $a_{11}$, $a_{00}$, $a_{{\textstyle\scriptscriptstyle-}1{\textstyle\scriptscriptstyle-}1}$. The same parameters enter the canonical amplitudes. For an on-shell decay, these parameters are constant; however, for $H\to VV$, they depend on the invariant mass of the off-shell $V$ boson $m_{V^{*}}$, or equivalently, on the modulus of the three-momenta in the Higgs rest frame, which we denote as $q$.
@@ 3 Density operators for Higgs decay
Notice that $\rho_{LS_{1}S_{2}}$ as calculated from this equation always corresponds to a pure state. As mentioned, $a_{11}$, $a_{00}$, $a_{{\textstyle\scriptscriptstyle-}1{\textstyle\scriptscriptstyle-}1}$ depend on $m_{V^{*}}$. The kinematical distribution of this quantity, calculated with Madgraph [26] at the leading order, is presented in Fig. 1 for $H\to WW$ and $H\to ZZ$. In practice, Eq. (8) can be used to calculate the density operator within some interval of $m_{V^{*}}$. For example, in $H\to ZZ$ with $m_{Z^{*}}\in[18,32]$ GeV (which amounts to 55% of the events) the exact density operator calculated with Monte Carlo describes a pure state up to corrections of the order of $0.5\%$. Likewise, for $H\to WW$ with $m_{W^{*}}\in[30,44]$ GeV (49% of the events) the density operator corresponds to a pure state up to corrections of 1%. Therefore, using (8) is quite a good approximation even for bin widths of the order of 10 GeV.
@@
In order to obtain $\rho_{LS_{1}S_{2}}$ for the full decay phase space we use Monte Carlo calculations of $gg\to H\to ZZ\to e^{+}e^{-}\mu^{+}\mu^{-}$ and $gg\to H\to W^{+}W^{-}\to\ell^{+}\nu\ell^{-}\nu$ with Madgraph, using $7\times 10^{6}$ and $10^{7}$ events, respectively. In $H\to ZZ$, we label the boson with largest invariant mass as $V_{1}$, and in $H\to W^{+}W^{-}$ we select $V_{1}=W^{+}$. (In any case, the predictions are symmetric under interchange $1\leftrightarrow 2$.) We divide the $m_{V^{*}}$ range in 2 GeV intervals and, within each bin ‘$k$’, the values of $a_{11}$, $a_{00}$ and $a_{{\textstyle\scriptscriptstyle-}1{\textstyle\scriptscriptstyle-}1}$ are extracted from Monte Carlo pseudo-data (see section 5 for details) using parton-level information. [^1] As a cross-check, we do the same in 5 GeV intervals. The numerical results are the same, with differences in the range $10^{-4}-10^{-3}$. The density operator $\rho_{LS_{1}S_{2}}^{(k)}$ for that bin is calculated using (8), and the theoretical prediction for $\rho_{LS_{1}S_{2}}$ in the full $m_{V^{*}}$ range is obtained by summing the operators $\rho_{LS_{1}S_{2}}^{(k)}$ in the different bins, with the appropriate weight. This density operator no longer describes a pure state, but it is quite close. For $H\to ZZ$, the principal eigenvector has eigenvalue 0.966 (where unity would correspond to a pure state), and in $H\to WW$ the principal eigenvector has eigenvalue 0.963.
@@ 4 Spin entanglement and Bell inequalities
with a sum over all indices. The coefficients satisfy the relations
@@ 6 Sensitivity in $H\to ZZ$
[…]
@@
The improvement of statistical uncertainties achieved with this binned reconstruction method is remarkable. If systematic uncertainties are under control, it allows to establish genuine tripartite entanglement in $H\to ZZ$ decays beyond the $5\sigma$ level at the current LHC run. Also, it allows to verify the violation of Bell inequalities between the two boson spins beyond the $5\sigma$ level.

### v2

@@ 1 Introduction
High-energy physics experiments offer a new playground to test quantum entanglement at the energy frontier, with proposals to test the spin entanglement of pairs of top quarks [1, 2, 3, 4, 5, 6, 7, 8], muons [9], $\tau$ leptons [10], weak bosons [11, 12, 13, 14, 15, 16, 17], and particles of different spin [18, 19, 20]. It is well known that, in addition to spin, particle production and decay involves orbital angular momentum (hereafter referred to as $L$). Therefore, in the aforementioned examples not only the spins of the two particles $S_{1}$, $S_{2}$ may be entangled, but they can also be entangled with $L$, as a consequence of total angular momentum conservation. An obvious step further is then to test the tripartite entanglement between $S_{1}$, $S_{2}$ and $L$ in these processes. The Hilbert space for $L$, in which the eigenstates $|lm\rangle$, with $l=0,\dots\infty$ and $m=-l,\dots,l$ form an orthonormal basis, has infinite dimension in general. However, in decay processes the conservation of total angular momentum implies that only a finite-dimensional subspace $\mathcal{H}_{L}$ is relevant. For example, in Higgs boson decays to weak boson pairs $VV$, $V=W,Z$, we have $l\leq 2$, and $\mathcal{H}_{L}$ is 9-dimensional. Likewise, in $H\to\tau^{+}\tau^{-}$, $\mathcal{H}_{L}$ has dimension four, with $l\leq 1$.
@@ 2 Higgs decay amplitudes
The angles $(\theta,\phi)$ correspond to the first decay product (with helicity $\lambda_{1}$). In the case of a scalar decay, $D^{0}_{00}=1$, and therefore the helicities are equal. Furthermore, for a scalar decay into vector bosons, the helicity amplitudes are parameterised by three quantities $a_{11}$, $a_{00}$, $a_{{\textstyle\scriptscriptstyle-}1{\textstyle\scriptscriptstyle-}1}$. The same parameters enter the canonical amplitudes. For an on-shell decay, these parameters are constant; however, for $H\to VV$ they depend on the invariant mass of the off-shell $V$ boson $m_{V^{*}}$, or equivalently on the modulus of the three-momenta in the Higgs rest frame, which we denote as $q$.
@@ 3 Density operators for Higgs decay
Notice that $\rho_{LS_{1}S_{2}}$ as calculated from this equation always corresponds to a pure state. As mentioned, $a_{11}$, $a_{00}$, $a_{{\textstyle\scriptscriptstyle-}1{\textstyle\scriptscriptstyle-}1}$ depend on $m_{V^{*}}$. The kinematical distribution of this quantity, calculated with Madgraph [26] at the leading order (LO), is presented in Fig. 1 for $H\to WW$ and $H\to ZZ$. In practice, Eq. (8) can be used to calculate the density operator within some interval of $m_{V^{*}}$. For example, in $H\to ZZ$ with $m_{Z^{*}}\in[18,32]$ GeV (which amounts to 55% of the events) the exact density operator calculated with Monte Carlo describes a pure state up to corrections of the order of $0.5\%$. Likewise, for $H\to WW$ with $m_{W^{*}}\in[30,44]$ GeV (49% of the events) the density operator corresponds to a pure state up to corrections of 1%. Therefore, using (8) is quite a good approximation even for bin widths of the order of 10 GeV.
@@
In order to obtain $\rho_{LS_{1}S_{2}}$ for the full decay phase space we use Monte Carlo calculations of $gg\to H\to ZZ\to e^{+}e^{-}\mu^{+}\mu^{-}$ and $gg\to H\to W^{+}W^{-}\to\ell^{+}\nu\ell^{-}\nu$ with Madgraph at LO, using $7\times 10^{6}$ and $10^{7}$ events, respectively. In $H\to ZZ$, we label the boson with largest invariant mass as $V_{1}$, and in $H\to W^{+}W^{-}$ we select $V_{1}=W^{+}$. (In any case, the predictions are symmetric under interchange $1\leftrightarrow 2$.) We divide the $m_{V^{*}}$ range in 2 GeV intervals and, within each bin ‘$k$’, the values of $a_{11}$, $a_{00}$ and $a_{{\textstyle\scriptscriptstyle-}1{\textstyle\scriptscriptstyle-}1}$ are extracted from Monte Carlo pseudo-data (see section 5 for details) using parton-level information. [^1] As a cross-check, we do the same in 5 GeV intervals. The numerical results are the same, with differences in the range $10^{-4}-10^{-3}$. The density operator $\rho_{LS_{1}S_{2}}^{(k)}$ for that bin is calculated using (8), and the theoretical prediction for $\rho_{LS_{1}S_{2}}$ in the full $m_{V^{*}}$ range is obtained by summing the operators $\rho_{LS_{1}S_{2}}^{(k)}$ in the different bins, with the appropriate weight. This density operator no longer describes a pure state, but it is quite close. For $H\to ZZ$, the principal eigenvector has eigenvalue 0.966 (where unity would correspond to a pure state), and in $H\to WW$ the principal eigenvector has eigenvalue 0.963.
@@ 4 Spin entanglement and Bell inequalities
with a sum over all indices. Because $\rho_{S_{1}S_{2}}$ is Hermitian, the coefficients satisfy the relations
@@ 6 Sensitivity in $H\to ZZ$
In this section we assess the statistical uncertainty in the measurement of various entanglement observables in $pp\to H\to ZZ\to 4\ell$ the LHC, using Run 2 + Run 3 data, and at the HL-LHC. For the calculation of the expected number of events we use state-of-the art values of the Higgs production cross section and branching ratio into four electrons or muons. The cross section at next-to-next-to-next-to-leading order is 48.61 pb, 52.23 pb and 54.67 pb at centre-of-mass energies of 13, 13.6 and 14 TeV [34], and the Higgs branching ratio decay into four leptons (electrons or muons) is $1.24\times 10^{-4}$ [34]. The assumed luminosities are 350 fb-1 for Runs 2+3 and 3 ab-1 for HL-LHC. In order to have a more realistic estimate of the number of events in each case a lepton detection efficiency of 0.7 is assumed, yielding an overall detection efficiency of 0.25. This efficiency accounts for the minimum transverse momentum ($p_{T}$) thresholds required for lepton detection. We do not include any trigger requirement. The presence of four leptons from the Higgs decay, some of them with significant $p_{T}$, is expected to fulfill one or many of the trigger conditions for one, two, or three leptons [35]. In addition, we include the efficiency of the invariant mass cut required to remove the interference in same-flavour final states (see the appendix). Overall, the expected number of events for Runs 2+3 and HL-LHC are $N=490$ and $N=4500$, respectively.
@@
The improvement of statistical uncertainties achieved with this binned reconstruction method, with respect to previous results for $S_{1}$-$S_{2}$ entanglement [12], is remarkable. If systematic uncertainties are under control, it allows to establish genuine tripartite entanglement in $H\to ZZ$ decays beyond the $5\sigma$ level at the current LHC run. Also, it allows to verify the violation of Bell inequalities between the two boson spins beyond the $5\sigma$ level.